In a triple riveted butt joint with two equal cover strips, having zigzag


arrangement of riveting the thickness of the plate 10 mm. Consider allowable


tensile, crushing and shear stress are 100 MPa, 150 MPa and 80 MPa


respectively. Calculate: - (i) diameter of rivet and (ii) pitch of rivet joint.

1
Expert's answer
2021-11-08T08:49:25-0500

diameter of rivet "d = 6 \sqrt t = 6\sqrt10= 18.97 \approx20" mm

pitch of rivet joint.

The shear strength of one rivet in double shear 

"p_{s1}= 1.75\times \frac{\pi}{4}d^2\times \tau_s"

The crushing strength of one rivet  

"P_{c1}=td\sigma_c"

The rivet is weaker in shearing. Equating tearing strength of plate with a shearing strength of 3 rivets in a pitch length,

"(p-d) t \sigma_t= 3\times 1.75\times \frac{\pi}{4}d^2\times \tau_s"

"p=\frac{3\times 1.75\times \frac{\pi}{4}20^2\times 80}{10\times150}+20=107.96mm"